Give the steps you will use to separate the variable and then solve the equation:
5m+7=17

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

First we have to subtract 7 to the both sides of the equation,
Then, we get,
= 5m + 7 - 7 = 17 - 7
= 5m = 10
Now,
We have to divide both sides of the equation by 5,
Then, we get,
= 5m/5 = 10/5
= m= 2

Thanks